A stream during late winter has a temperature of 5 °C and 5 mg DO/L.

a) What is the percent saturation of DO? There is a spell of hot weather and the water temperature increases to 15 °C (at the same 5 mg DO/L).

b) What is the new percent DO saturation?

c) Even though the DO concentration was the same in terms of mg/L, why do these percent saturation values differ?

d) What does this mean qualitatively for aquatic and fish health under the two scenarios? Assume atmospheric pressure of 1 bar.
Answer #1

Answer a) At 5°C saturation DO = 12.75 mg/l thus percent saturation of DO = (5/12.75) *100 = 39.21 %

Answer b) At 15°C saturation DO = 10.07 mg/l thus percent saturation = (5/10.07)*100 = 49.65% = new saturation

Answer c) The values of percentage saturation differs because of the change in saturated dissolved oxygen capacity of water.

Answer d) Under the two scenarios the second scenario is better as percentage saturation is more. Thus fish and aquatic life would have better health if 5mg/l of dissolved oxygen is present at 15°C.